A person is swinging around a horizontal circle at the end of a rope on a fairground ride. The person weighs 86 kg, at the radius of the circle is 5.3 m. The rope is taut and horizontal as the person swings around at a constant speed of 7.9 m/s

What is the tension force in the rope in Newtons? Give your answer to the nearest whole number?
1
Expert's answer
2019-11-26T10:43:49-0500

Here V= 7.9 m/sec , Mass , M = 86 kg , Radius , R= 5.3 m


Tension Force T will be equal to the centripetal force .


So , The Tension T ="\dfrac{MV^2}{R}" = "\dfrac{86\times 7.9 \times 7.9}{5.3}" = 1013 Newton